# Knowledge-aware Parsimony Learning: A Perspective from Relational Graphs

### v2

@@ 8 Conclusion
This paper introduces an alternative way to develop next-generation learning techniques instead of scaling law. By leveraging the duality between data and knowledge, our method extracts symbolic logic and physical laws during the learning process and applies combinatorial generalization to various tasks. This approach effectively overcomes the limitations of traditional scaling methods. Experimental results demonstrate that our framework significantly improves model performance, showcasing its ability to achieve parsimony on model, training and interpretation. These findings underscore the potential of integrating knowledge into machine learning models, offering a promising direction for future research and applications.

### v3

@@ 8 Conclusion
This paper introduces an alternative way to develop next-generation learning techniques instead of brute-force scaleup. By leveraging the duality between data and knowledge, our method extracts symbolic logic and physical laws during the learning process and applies combinatorial generalization to various tasks. This approach effectively overcomes the limitations of traditional scaling methods. Experimental results demonstrate that our framework significantly improves model performance, showcasing its ability to achieve parsimony on model, training and interpretation. These findings underscore the potential of integrating knowledge into machine learning models, offering a promising direction for future research and applications.
Acknowledgment
Q. Yao work is supported by National Key Research and Development Program of China (under Grant No.2023YFB2903904), National Natural Science Foundation of China (under Grant No.92270106) and Beijing Natural Science Foundation (under Grant No.4242039). J. Kwok is supported in part by the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16202523 and C7004-22G-1). Finally, we thanks suggestive comments from reviewers, Wenguang Chen and Xuefei Ning for helping us improve the quality of this perspective.
Conflict of Interest
The authors have no conflicts of interest to report.